OpenStudy (raffle_snaffle):

How many molecules (not moles) of NH3 are produced from 1.77×10−4 g of H2?

OpenStudy (raffle_snaffle):

@Photon336

OpenStudy (raffle_snaffle):

@Preetha

OpenStudy (raffle_snaffle):

6.25x10^18 molecules but it isn't right

OpenStudy (jebonna):

Have you been given an equation with this question?

OpenStudy (raffle_snaffle):

no, are we suppose to use pv=nrt?

Vocaloid (vocaloid):

I'm taking a stab at what the chemical equation might be N2 + 3 H2 ---> 2 NH3 ?

Vocaloid (vocaloid):

it's a bit unclear with the info given though

OpenStudy (raffle_snaffle):

Well that is what's given. So I was suppose to balance the equation and now what do I need to do?

OpenStudy (raffle_snaffle):

@Photon336

OpenStudy (cuanchi):

1. convert the g of H2 to moles of H2 1.77×10−4 g of H2 x 1 mol H2/ 2 g H2 = xxx mole H2 2. by the stoichiometry of the reaction you see 2 mol NH3 / 3 mol H2 xxx mole H2 x 2 mol NH3 / 3 mol H2 = yyyy mole NH3 3 convert mol of NH3 to molecles of NH3 yyyy mole of NH3 x 6.02 x 10^23 molecules of NH3/ 1 mol NH3 = zzzz molecules of NH3 ------------------------------------------- in one equation with 3 conversion factors: 1.77 x 10-4 g x (2 molH2/2gH2) x (2molNH3 /3 mol H2) x ( 6.02x10^23 molec NH3/ 1 molNH3) = zzz molec of NH3